1) 2000 because it is the closest number to 4000
2) 700*23/100 = 161
3) (-6+4)/(6-1) = -2/

Found 6 solutions by usyim88hk, tinogonzales, JimMarshall, Robee16, The Troll Kid, ikleyn:
Answer by usyim88hk(158)

1) 2000 because it is the closest number to 4000
2) 700*23/100 = 161
3) (-6+4)/(6-1) = -2/5

Answer by tinogonzales(1)

1) 2000 because it is the closest number to 4000
2) 700*23/100 = 161
3) (-6+4)/(6-1) = -2/

Answer by JimMarshall(82)

2000 because it is the closest number to 4000
2) 700*23/100 = 161
3) (-6+4)/(6-1) = -2/5 Ans

Answer by Robee16(18)
Answer by The Troll Kid(22)
Answer by ikleyn(52780)

.

    N O N S E N S E